OpenStudy (anonymous):

A radioactive isotope has a half-life of one week. How much of it will remain after one month? Show all calculations leading to an answer OR explain your answer with 3 – 4 complete sentences.

OpenStudy (anonymous):

@hugekabab

OpenStudy (anonymous):

ooo half life ... its been a while since i did these ... let me think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok listen closely

OpenStudy (anonymous):

assume that the isotope starts at x units

OpenStudy (anonymous):

. after 1 week half will be gone, or x/2, after the second week half of what remains will be gone. x/2 /2 or x/4, after three weeks x/8 and after 4 weeks, x/16 or 1/16 of the starting amount

OpenStudy (anonymous):

get it ?

OpenStudy (anonymous):

yes i kinda already understood half life basically each weach this isotope will loose half of it total life left

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

btw i gotta ask ... is this for hig school ?

OpenStudy (anonymous):

so W1=1/2 W4=1/16

OpenStudy (anonymous):

so at 1 month it will have 1/16 of its life

OpenStudy (anonymous):

thats what i think

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

half of the orignal in the first week ... half of the half left in the second week (quater of the orignal) and so on

OpenStudy (anonymous):

get the sequence ?

OpenStudy (anonymous):

and just to clarify on the last question we did i divided those numbers and got 60041537135.7 and i add KG-J

OpenStudy (anonymous):

yes i got it thank you so much

OpenStudy (anonymous):

just kg .... cause its just simple mass ... nothing more ..and btw ... do the math again just to be sure yu have the figures right ..

OpenStudy (anonymous):

i was kinda hoping for a medal ...its like my first time here on open study ... and damn is this place cool

OpenStudy (anonymous):

i literally divided the two as is

OpenStudy (anonymous):

yeah yu did multiply it by C right ? the speed of light

OpenStudy (anonymous):

yu did right ... as long as you multiplied C as well

OpenStudy (anonymous):

no i divided m=e/c^2

OpenStudy (anonymous):

yeah divide ...sorry my bad

OpenStudy (anonymous):

its correct thrn

OpenStudy (anonymous):

then *

OpenStudy (anonymous):

C must be in m/s

OpenStudy (anonymous):

and what about the ^2 over c? is that encompassed in the number i used for the speed of light

OpenStudy (anonymous):

yes .. thats the speed of light squared

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well thank you than

OpenStudy (anonymous):

no problem ... always happy to help ..